# Robust Radar Mounting Angle Estimation in Operational Driving Conditions

### v1

@@ IV-D Convergence
The previous results demonstrated that the proposed method achieves accurate and robust radar mounting angle estimation. However, those results were either evaluated per scene (Figure 5) or averaged across all scenes (Table II). For realistic driving scenarios, it is equally important to assess how quickly (in terms of seconds) the estimation converges to a reliable and stable value. To examine this, Figure 6 presents the mean absolute error (MAE) and error variance as functions of the time interval, i.e., the number of frames processed by each method. For MAE (Figure 6a), the proposed method achieves stable and low errors even with only a few seconds, whereas the two reference methods require much more time before producing accurate results. For error variance (Figure 6b), the proposed method consistently benefits from longer time intervals, yielding smaller variances. In contrast, the baseline methods show no clear convergence: their variance decreases initially, but then increases and fluctuates strongly with longer intervals.
Finally, it is worth noting that these results are based on the 64 test scenes from Radar 3, the forward-looking radar with many moving objects in its field of view. The result demonstrates that the proposed method can produce accurate and confident mounting angle estimates even under challenging conditions, within a short period of time.
@@ IV-F Further Exploration
To better understand the performance trade-offs, the proposed method was modified accordingly while still using the frame weights from Equation (13). Results are shown in Table IV. Compared with Table II, the performance gap between different formulations is now much smaller. Since the main difference between these methods and the baselines from the literature lies in the radar motion and frame weighting, this suggests that the key bottleneck in robust and accurate mounting angle estimation is the high proportion of outliers and sparse radar measurements. For the chosen dataset, the weighted mean solution appears most practical: it uses the simplest model, runs the fastest, and provides performance comparable to more complex techniques. If the IMU scale factor deviates significantly from 1, the weighted LSQ should instead be used. Finally, if the radar has limited resolution in azimuth and radial velocity, weighted ODR may yield better results.
@@ V Conclusions
This paper presented a novel signal processing pipeline for robust estimation of radar mounting angles under operational driving conditions. Accurate external calibration of automotive radars, and in particular their mounting angles, is crucial for the safe operation of autonomous vehicles. To address this problem, we proposed an odometry-based approach that combines a neural network radar motion estimator with an IMU measurement model for de-biasing and scale factor estimation. The mounting angle and IMU scale factor are then jointly estimated using a wLSQ formulation based on a Taylor-series linearization. The proposed pipeline was validated on the challenging RadarScenes dataset. Results demonstrate both robustness and accuracy: the method outperforms existing approaches, achieves consistent estimates across diverse and realistic driving scenarios, and avoids the constraints of controlled environments, specially designed radar targets, or tailored driving routes. Importantly, while the formulation involves a single linearization step, in practice this single step was sufficient to achieve convergence, requiring no further iteration. For future work, extending the framework to sensor fusion represents a promising direction. Combining multiple radar sensors or integrating complementary modalities (e.g., cameras or lidars) may further mitigate the impact of dynamic objects in the field of view and enhance calibration accuracy under complex real-world conditions.

### v2

@@ IV-D Convergence
The previous results demonstrated that the proposed method achieves accurate radar mounting angle estimation with low scene-to-scene variability. However, those results were either evaluated per scene (Figure 5) or averaged across all scenes (Table II). For realistic driving scenarios, it is equally important to assess how quickly (in terms of seconds) the estimation converges to a stable value.
To examine this, Figure 6 presents the mean absolute error (MAE) and error variance as functions of the time interval, i.e., the number of frames processed by each method. For MAE (Figure 6a), the proposed method achieves low estimation error within only a few seconds of driving, whereas the two reference methods require much more time before producing comparable accuracy. For error variance (Figure 6b), the proposed method consistently benefits from longer time intervals, yielding smaller variances. In contrast, the baseline methods show no clear convergence: their variance decreases initially, but then increases and fluctuates strongly with longer intervals.
Finally, it is worth noting that these results are based on the 64 test scenes from Radar 3 of the RadarScenes dataset, the forward-looking radar with many moving objects in its field of view. The result demonstrates that the proposed method can produce accurate mounting angle estimates with low variance within a short period of driving, even under challenging traffic conditions and with variation in velocity and trajectories of the ego-vehicle.
@@ IV-F Further Exploration
To better understand the performance trade-offs, the proposed method was modified accordingly while still using the frame weights from Equation (13). Results are shown in Table IV. Compared with Table II, the performance gap between different formulations is now much smaller. Since the main difference between these methods and the baselines from the literature lies in the radar motion and frame weighting, this suggests that a major limiting factor in practical mounting angle estimation is the high proportion of outliers and sparse radar measurements. For the chosen dataset, the weighted mean solution appears most practical: it uses the simplest model, runs the fastest, and provides performance comparable to more complex techniques. If the IMU scale factor deviates significantly from 1, the weighted LSQ should instead be used. Finally, if the radar has limited resolution in azimuth and radial velocity, weighted ODR may yield better results.
@@ V Conclusions
This paper presented a novel signal processing pipeline to address the problem of estimating radar mounting angles under operational driving conditions. Accurate external calibration of automotive radars, and in particular their mounting angles, is crucial for the safe operation of autonomous vehicles. To address this problem, an odometry-based approach was proposed that combines a neural network-based radar motion estimator with an IMU measurement model for bias and scale factor compensation. The mounting angle and IMU scale factor are then jointly estimated using a wLSQ formulation based on a Taylor-series linearization. The proposed pipeline was validated on the challenging RadarScenes dataset, which includes diverse traffic scenes as well as velocity and trajectory variations of the ego-vehicle.
Experimental results demonstrate that the method achieves accurate mounting angle estimates with low variability across diverse and realistic driving scenarios, while avoiding the need for controlled environments, specially designed radar targets, or tailored driving routes. In addition, the estimation converges within a short period of driving time. Although the formulation involves a single linearization step, experimental results show that this is sufficient in practice, requiring no further iteration. For future work, extending the framework to sensor fusion represents a promising direction. Combining multiple radar sensors or integrating complementary modalities (e.g., cameras or lidars) may further mitigate the impact of dynamic objects in the field of view and enhance calibration accuracy under complex real-world conditions.